# Tunable Topological Phases in Quantum Kirigamis

### v1

@@ Abstract
Introduction: The possibility of mechanically deforming mesoscopic quantum devices holds promise for the next generation of quantum technologies Rogers et al. (2010); Blees et al. (2015); Ouyang et al. (2023); Liu et al. (2018); Jin and Yang and also raises interesting theoretical questions Castro et al. (2018); Flachi and Vitagliano (2019); Imaki and Yamamoto (2019). The experimental demonstration of graphene Kirigamis Blees et al. (2015), their bending rules Grosso and Mele (2015), and high throughput computational search Hanakata et al. (2018) have established them as a unique platform. While a lot of work has gone into engineering the mechanical properties of such systems Liu et al. (2022); Jin and Yang , little effort has been given to uncover the transport properties of such systems Bahamon et al. (2016); Mortazavi et al. (2017).
@@
In this work we pose the question, can the mechanical modulation engineer a quantum phase transition in such Kirigami-based systems? In particular, in recent years, topological phases of matter have ushered in a new paradigm within condensed matter research where materials can host robust quantized transport on the edges even while the bulk remains insulating. Such systems are characterized by topological invariants and are protected by discrete symmetries Ludwig (2016); Chiu et al. (2016); Hasan and Kane (2010); Qi and Zhang (2011); Asbóth et al. (2016). We explore whether Kirigami networks built out of such topological materials can show topological phase transitions as a function of the mechanical deformation of the structure. We establish theoretically that indeed such phase transitions are realizable in these systems as a function of deformation parameters. Using a multi-scale approach, starting from minimal models, effective junction modelling, and scattering-matrix calculations we show that such systems can be effectively modelled as ‘tunable’ Chalker-Coddington networks Chalker and Coddington (1988); Ho and Chalker (1996); Kramer et al. (2005); Lee et al. (1993); Snyman et al. (2008); Cho and Fisher (1997), usually studied for disordered topological systems Charles et al. (2020); Potter et al. (2020); Obuse et al. (2012); Hu et al. (2015); Chalker et al. (2001); De Beule et al. (2021) where the deformation of the network can be included to determine the effective parameters of the network itself.
@@
We model a foldable Kirigami as a network of two-dimensional lattice blocks attached at the vertices as shown in Fig. 1(a). If each of the blocks themselves hosts a topological phase such that it retains a chiral edge state, then the vertices can be effectively modeled by a scattering matrix characterized by reflection and transmission amplitudes $r$ and $t$ (see the zoomed region in Fig. 1(a)). Interestingly as the structure unfolds, the angle $2\theta$ between the blocks can be tuned as shown in Fig. 1(b). Then the natural question arises: given any single block containing topological or trivial phases as a function of any microscopic parameter $M$ (say); can the mechanical deformation $\theta$ engineer a similar phase transition? As we will show that indeed $\theta$ can effectively engineer a phase transition where the effective Hamiltonian for the network shows Dirac cone closings as shown schematically in Fig. 1(d), rendering the network to become trivial even when each of the blocks are themselves topological. The complete phase diagram is shown in Fig. 1(c).
@@
Model: We model the Kirigami with a periodic arrangement of blocks with associated nodes governed by scattering matrices. We consider the microscopic Hamiltonian governing each of the blocks to the paradigmatic Bernevig-Hughes-Zhang (BHZ) model Bernevig et al. (2006). The Hamiltonian in momentum space is given by
@@
where $M$ is the parameter which tunes topological and trivial phases such that for $M>0$ and $M<-4$ we have a trivial phase, while for $20>M>-$ 2 and $-2>M>-4$ we have Chern insulating phases with Chern number 1 and -1 respectively. All parameters are measured in terms of the bare hopping scale which is set to unity. All through this work we will assume that the lattice constant $a$ is set to unity and is much smaller than the typical size of the block such that each block by itself can be considered a thermodynamically large system. When $M$ is such that any block is in a topological phase the bulk of the each block remains gapped while the low-energy excitations are chiral and are present only on the boundaries. In particular Chern number $C=-1$ ($C=1$) phase leads to clockwise (anticlockwise) edge states. Thus the junction between adjoining blocks are governed by a scattering matrix between the edge states, which we discuss next.
Any node in this network is composed of two incoming ($1,3$) and two outgoing channels ($2,4$) (see Fig. 2(a) inset). Thus the scattering matrix $S$ is defined by two parameters $r$ and $t$ which are in general $\theta$ dependent, such that
@@
where $|r|^{2}+|t|^{2}=1$ Kramer et al. (2005). In general $r,t$ are complex amplitudes, which we will choose to be real numbers and discuss later on the effect of phases. To estimate the functional behavior of $r,t$ on $\theta$ we use multiple methods discussed next.
Figure 2: Junction physics: (a) Schematic showing that an incident electronic wave packet (from edge channel 1) on a node, gets reflected (on channel 2) and transmitted (on channel 3) by scattering matrix elements $r$ and $t$ respectively. $\lambda(\theta)$ is inter-edge hopping amplitude between sites separated by a distance $d(\theta)$ within a width $w$. (b) Behavior of $r(\theta),t(\theta)$ as a function of $\theta$ using wave-packet dynamics (see text) for $M=-3.6,\alpha=1,\beta=1,w=5,N=1250$. Here $N$ is the total number of sites in both the blocks forming the junction. (c) Schematics of a one-dimensional non-Hermitian system which also models the node and scattering physics. In (I) similar inter-edge coupling is introduced as in (a) to obtain $r(\theta),t(\theta)$ while in (II) the node itself is replaced by four sites with intersite non-Hermitian hoppings $r(\theta)$ and $t(\theta)$. (d) Behavior of $r(\theta)$ and $t(\theta)$ with $\theta$ using wave-packet dynamics on non-Hermitian system (I) ($\alpha=1.5,\beta=0.25,w=5$). Using these obtained $r,t$ in the model (II), reproduces the same curve (not shown).
Wave-packet dynamics: A state localized on a single site on the edge is initially projected in the low energy subspace $|E|<\Delta_{g}$ where $\Delta_{g}$ is the bulk gap in the topological regime. This leads to a finite projection into the chiral edge state which moves unidirectionally under time evolution (see schematic in Fig. 2(a), and Supplemental Material (SM) sup for details). After the packet leaves the junction the individual probability densities on the edges (2,4) given by $p_{2}$ and $p_{4}$ are evaluated numerically. This leads to $r\propto\sqrt{p_{2}}$ and $t\propto\sqrt{p_{4}}$ with a normalization constraint $r^{2}+t^{2}=1$. In the microscopic dynamics, in order to realistically model a smooth junction we introduce weak inter-block hoppings where fermions can hop from edges facing each other by an exponentially falling hopping amplitude $\lambda(\theta)=\alpha\exp(-\beta d(\theta))$ where $\alpha,\beta$ allows us to modulate the coupling between edges and $d(\theta)$ is the distance between two sites on the facing edges within a finite range $w$ of the Kirigami junction (see Fig. 2(a)). Here we fix $\alpha,\beta=1,w=5$ unless mentioned otherwise. Choosing $M=-3$ which is deep in the topological phase with Chern number $C=-1$ (clockwise edge states) we evaluate $r,t$ for all values of $\theta$. We find an intriguing dependence, where $t(\theta)\sim 1$ near both $\theta\sim 0$ and $\sim\frac{\pi}{2}$ while $r(\theta)>t(\theta)$ near $\theta\sim\frac{\pi}{4}$ (see Fig. 2(b)). We have checked that this behavior is independent of the system size (see SM sup ). We note that the exact values depend on the choice of $\lambda(\theta),w$, and $M$, however, the qualitative behavior of these coefficients do not sup . This motivates us to present an effective one-dimensional edge model to qualitatively capture the physics of this junction.
Effective non-Hermitian chain: In order to capture the effective scattering of the one-dimensional edge mode we model the chiral edge state with a one-dimensional Hatano-Nelson model Hatano and Nelson (1996, 1997) where the fermions are allowed to hop in just one direction. This uses the mapping of boundary theories of $d$ dimensional topological phases to $d-1$ dimensional non-Hermitian systems Lee et al. (2019); Lieu (2018); Shen et al. (2018); Schindler et al. (2023); more physically such a scattering mapping prevents any backscattering from the junction as is true for chiral systems. To make this connection more concrete, we consider the case where we wish to model the effective Hamiltonian of just a chiral edge state: $H_{\text{edge}}\sim v_{F}kc^{\dagger}_{k}c_{k}$ where $v_{F}$ is the edge state Fermi velocity. A non-Hermitian Hatano-Nelson (HN) Hamiltonian of the form
$H_{\text{edge}}=\sum_{i}-c^{\dagger}_{i+1}c_{i}$ (3)
where $i$ runs over the one-dimensional chain, stabilizes a unidirectional finite current at finite densities in a steady state mimicking a chiral edge state Hatano and Nelson (1997); Lee et al. (2019); Banerjee et al. (2022). We thus model the junction using an effective non-Hermitian system. We use four one-dimensional HN chains to form a junction with non-reciprocal hoppings in the direction of the edge currents as shown in Fig. 2(c, I). Similar to the BHZ system, interchain couplings are added in form of $\lambda(\theta)$. Again doing a wavepacket evolution we calculate $r(\theta),t(\theta)$ which is shown in Fig. 2(d). The qualitative similarity between (b) and (d) shows that the non-Hermtian scattering junction is indeed able to capture the scattering physics of the microscopic BHZ system. Interestingly, $r(\theta),t(\theta)$ can themselves be interpreted as a parameter to define another non-Hermitian junction as shown in Fig. 2(c, II) which subsumes the other particulars of the junction leading to identical reflection and transmission amplitudes. These results, taken together, illustrate that the scattering amplitudes are indeed dependent on the angle of the junction between each of the Kirigami blocks independent of microscopic parameters or modelling protocols.
Chalker Coddington networks: Having defined the property of every junction we discuss the phases realized due to the complete Kirigami network. The effective hopping problem due to the transfer matrix can be designed as a unitary evolution where an initial state evolves as Charles et al. (2020)
$|\Psi_{f}\rangle=U_{\text{eff}}|\Psi_{i}\rangle$ (4)
where the effective evolution operator, $U_{\text{eff}}$ is decided by the scattering matrix structure of the network (see SM sup ). The $U_{\text{eff}}$ is constructed using a median lattice where each edge of the Kirigami block is replaced by a site such that the system has a 16-site unit cell. The ‘hopping’ elements are represented by $r(\theta)$ and $t(\theta)$ with appropriate convention depending on the direction of the edge current. Given the network is translationally invariant, in order to see if it retains any edge states we set it up on a ribbon geometry which is finite in the $y$ direction and periodic in the $x$ direction. We find the eigenvalues of $U_{\text{eff}}$ which are of the form $e^{-i\epsilon}$ and plot $\epsilon$ as a function of $k_{x}$. Note that $r,t$ enter as parameters in the transfer matrix.
When $r<t$ and $\theta<\theta_{1c}$, we find linearly dispersing edge states on the boundary (see Fig. 3(d)) reflecting that the topological character of the individual Kirigami block also leads to an overall chiral edge in the system (Fig. 3(a)). The edge state profile is shown in SM sup . However now as $\theta$ is increased, or the structure opens up, the individual scattering elements ($r(\theta),t(\theta)$) modulate approaching a critical point $\theta_{1c}$ where $r=t$. At this point, the system undergoes a phase transition where the $\epsilon$ shows a Dirac-like gap closing. After $\theta>\theta_{1c}$ the gap opens up again (Fig. 3(e)) albeit with no chiral mode on the complete network. This reflects that the Kirigami network has become a “trivial” insulator (see Fig. 3(b)). This is an example of a mechanical deformation-induced topological phase transition. With further increase in $\theta$ another critical point $\theta_{2c}$ is approached where the system transits back from the trivial to topological phase (see Fig. 3(c) and (f)). The specific values of $\theta_{1c}$ and $\theta_{2c}$ are dependent on both the microscopic parameters of the Hamiltonian governing each block, as well as the nature of the network. For instance in the case of the square network shown in Fig. 1(a), $\theta_{1c}$ and $\theta_{2c}$ are symmetric about $\pi/4$. For every value of $M$ numerically evaluated $r=t$ points leads to the phase diagram as shown in Fig. 1(c),(d). While we have chosen $r,t$ to be real, a uniform phase does not change $\theta_{1c},\theta_{2c}$. However relative scattering phase either uniform, or randomly varying from block to block can lead to both changes in critical angles and new phases Charles et al. (2020) which we do not explore here but is an interesting future direction.
We note that this analysis and formalism is only applicable when each of the Kirigami blocks resides in the topological regime i.e. $-4<M<0$. For other values of $M$ when each of the blocks is itself an insulator - mechanical deformation doesn’t render it topological. Within $-4<M<0$ each block has a topological phase transition at $M=0,-2$, and $-4$ where the system contains Dirac cones. Here again, the blocks are themselves metallic and this analysis breaks down. Given the $M=-2$ point sits between the two topological Chern insulator phases of Chern number ±$\pm 1$, $r=t$ by symmetry reasons. Overall the phase diagram is further symmetric about $M=-2$ since apart from the direction of the chiral edge states the physics remains the same.
It is interesting to point out the physics of the $r=t$ line in this system, particularly near $\theta\sim 0$. Given when each Kirigami block is in the topological regime - there exists an edge state with an edge localization length $\xi$, which is dependent on the bulk gap which in turn depends on $M$. Thus whether an edge wavepacket can transmit or reflect from the junction depends on whether $\xi\ll w$ or $\xi\gg w$ respectively. Therefore given a choice of $w$, there exists a critical value of $M$ even within the topological phase, defined as $M^{*}$ at which $\xi(M)\sim w$, therefore leading to a point where $r=t$. This value of $M^{*}$ is thus independent of the system size of the Kirigami block but only dependent on the value of $\theta$ and the parameters defining the junction. We further note tuning $\alpha,\beta$ which defines the strength of the junction can further tune the values of $r,t$ (see SM sup ).
In order to test this physics in another network, we set up the same problem however now on a Kirigami network where the network deforms between the triangular lattice to a Kagome-like network as a function of deformation angle $\theta$ (see Fig. 4(a),(b)). The microscopic Kirigami block is still assumed to be governed by a square lattice BHZ model. Given a different network structure the $U_{\text{eff}}$ changes. However again as a function of $\theta$ the system undergoes topological to trivial to topological phase transitions as shown in (see Fig. 4(c)-(e). What is interesting however is that the critical points $\theta_{1c}$ and $\theta_{2c}$ are now symmetric about $\theta=\frac{\pi}{3}$. Moreover the topological transition occurs at $r(\theta)=\frac{1}{2},t(\theta)=\frac{\sqrt{3}}{2}$. Thus the transitions can be found when $\tilde{r}(\theta)=\tilde{t}(\theta)$ where $\tilde{r}(\theta)\equiv 2r(\theta)$ and $\tilde{t}(\theta)\equiv\frac{2}{\sqrt{3}}t(\theta)$. The behavior of $\tilde{r}(\theta)$ and $\tilde{t}(\theta)$ for $M=-3.5$ is shown in Fig. 4(i). The spectrum when evaluated on a ribbon geometry clearly shows the edge states in the topological regimes (see Fig. 4(f-h)). This reflects the generality of this physics to various Kirigami systems and that critical angles for transitions can themselves be tuned by nature of the Kirigami structure.
Outlook: Metamaterials and mechanically engineered systems herald the new era of technological devices. A host of electronic systems based on such structures have been realized. However, the interplay of quantum phases and their transitions vis-a-vis their mechanical tunability has been little explored. In this work, we show that topological phases, in particular a Chern insulator which contains chiral edge channels can be tuned to a trivial phase by such mechanical tuning. Recent experimental progress, in particular, such Kirigami-based systems Jin and Yang have significantly expanded the capabilities to realise this physics. Our study places this idea on a concrete footing in the context of topological phases uncovering their mechanism in a simple setting. The effect of realistic strains, gradients in deformations, strong correlations, and interactions are some of the many future directions this work opens up. Equivalent physics in alternate platforms such as photonics and phononic systems may be another interesting direction to pursue.
Acknowledgement: We acknowledge fruitful discussions with G. Murthy, G. Sreejith, Diptarka Das, Arijit Kundu, Subrata Pachhal, Sudipta Dubey, Ritajit Kundu, Saikat Mondal and Soumya Sur. RS acknowledges funding from IIT Kanpur Institute Fellowship. AA acknowledges support from IITK Initiation Grant (IITK/PHY/2022010).
Supplemental Material for “Tunable Topological Phases in Quantum Kirigamis”
In this SM we discuss the numerical procedure to estimate $r(\theta),t(\theta)$ and present additional results on their dependence on various model parameters.
I Wave packet evolution and estimation of $r(\theta),t(\theta)$

### v2

@@ I Introduction
The possibility of mechanically deforming mesoscopic quantum devices holds promise for the next generation of quantum technologies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and raises interesting theoretical questions [11, 12, 13]. The experimental demonstration of graphene Kirigamis [7], their bending rules [14], and high throughput computational search [15, 16, 17] have established them as a unique platform. While a lot of work has gone into engineering the mechanical properties of such systems [18, 10], little effort has been given to uncover the transport properties of such systems [16, 19, 20, 21].
@@
In this work, we pose the question, can mechanical modulation engineer a quantum phase transition in such Kirigami-based systems? In particular, in recent years, topological phases of matter have ushered in a new paradigm where materials can host robust quantized transport on the edges even while the bulk remains insulating [22, 23, 24, 25, 26]. We explore whether Kirigami networks built out of such topological materials can show phase transitions as a function of the mechanical deformation. We establish theoretically that indeed such phase transitions are realizable in these systems. Using a multi-scale approach, starting from minimal models, effective junction modelling, and scattering-matrix calculations we show that such systems behave as ‘tunable’ Chalker-Coddington networks [27, 28, 29, 30, 31, 32], usually studied for disordered topological systems [33, 34, 35, 36, 37, 38, 39] where the deformation of the network can be included to determine the junction parameters of the network.
@@
We model a foldable Kirigami as a network of two-dimensional lattice blocks attached at the vertices (see Fig. 1(a)). If each of the block is topological such that it hosts a chiral edge state, then the vertices can be modeled by a scattering matrix characterized by reflection and transmission amplitudes $r$ and $t$ (see the zoomed region in Fig. 1(a)). Interestingly as the structure unfolds, the angle $2\theta$ between the blocks can be tuned as shown in Fig. 1(b). In the regime where a single block becomes topological as a function of a microscopic parameter $M$ (say), we show that $\theta$ can be used to engineer a topological phase transition where the effective Hamiltonian for the network shows Dirac cone closings as shown schematically in Fig. 1(d). This renders the network trivial even when each of the blocks are themselves topological. The complete phase diagram is shown in Fig. 1(c).
In section II and III we discuss the microscopic model and the physical properties of the Kirigami. In section IV we discuss the utility of the three-step methodology which we employ. In section V we present all the results and discuss the phase diagram of the system. In section VI we study another triangular Kirigami network using the same formalism. In section VII we conclude our work. Additional numerical results and parameter dependencies have been relegated to Appendix IX.
@@
Figure 2: Multiscale approach:  The problem is analyzed within a three-stage process where in the first step (a) wave-packet dynamics (WPD) is used to characterize a single node of junctions between Kirigami blocks (see section V.1). (b) An effective non-Hermitian junction (ENHJ) to model the complete network with two types of nodes $N_{1}$ and $N_{2}$ as shown with yellow and blue colors respectively (see section V.2). (c) A Chalker-Coddington network where a median lattice is used to understand the topology of the full network (see section V.3).
II Model
We model the Kirigami with a periodic arrangement of blocks with associated nodes governed by scattering matrices. We consider the microscopic Hamiltonian governing each of the blocks to be the paradigmatic Bernevig-Hughes-Zhang (BHZ) model [40]. The Hamiltonian in momentum space is given by
@@ II Model
where $M$ is the parameter that tunes topological and trivial phases such that for $M>0$ and $M<-4$ we have a trivial phase, while for $20>M>-$ 2 and $-2>M>-4$ we have Chern insulating phases with Chern number 1 and -1 respectively. All parameters are measured in terms of the bare hopping scale which is set to unity. We assume that every block is thermodynamically large while the lattice constant $a=1$ and is much smaller than the block size $L^{2}$. In the topological phase, the edge retains chiral edge states while the bulk states remain gapped. In particular Chern number $C=-1$ ($C=1$) phase leads to clockwise (anticlockwise) edge states. The whole network containing $L_{B}\times L_{B}$ number of Kirigami blocks can thus be represented by a tight-binding Hamiltonian with a total size of single-particle Hilbert space dimension $N_{T}=2L^{2}L^{2}_{B}$ since every site contains two orbitals $A$ and $B$. To understand the transport property of such a network we use a multi-scale methodology with three different stages applied to the morphology of the Kirigami structure. We first briefly discuss the Kirigamis we intend to study before detailing its transport properties.
III Kirigami morphology
Before we specifically comment on methods to calculate transport quantities it is useful to discuss the mechanical properties of Kirigamis and how they behave under deformation thus affecting the underlying microscopic Hamiltonian. Kirigamis, particularly of experimental and technological interests can either be parallel cut or cross cut [41, 42, 43]. The parallel-cut version of Kirigami drew a lot of theoretical interest due to its realization in graphene [7]. When strained, this goes through four stages of deformation [17, 16], of which the first two are primarily in plane while in the latter two, the structure twists out of the plane and starts tearing apart. In the first stage, even while the inter-atomic bond lengths do not change significantly the Kirigami itself expands due to structural cuts. In the second stage, the inter-atomic bonds themselves start expanding and bending which requires the evaluation of the effective fermionic hoppings dynamically [16]. On the other hand, cross-cut Kirigamis (also called auxetic structures [41, 42, 43, 44, 45]) have a larger window of in-plane deformations which are further tunable by controlling the thickness of Kirigami and the size of cuts. In our work, we will work with auxetic Kirigamis and assume the hopping parameters to be static such that all deformations are in the plane. Thus for an undeformed structure ($\theta=0^{\circ}$) of total length $l=LL_{B}$, at any angle $\theta$ the total length of the network changes to $l^{\prime}$ such that [44]
$l^{\prime}=l[\cos(\theta)+\sin(\theta)]$ (2)
where the effective strain can thus be estimated as $\epsilon_{s}=(l^{\prime}-l)/l$. It is known from experimental studies on auxetic metamaterials [42] that depending on the ratio of the thickness and junction width, the system can show in-plane deformations even for significant strains. While the direct translation of these mechanical properties into nanomaterials is far from obvious, approximately $30\%$ strain, corresponds to $\theta\sim 22^{\circ}$ (see eqn. (2)). In graphene systems [7] a strain of $\sim 240\%$ was realized, however, this necessarily introduces significant inhomogeneities and twisting out of the plane. In most of our study, we will assume homogeneous deformations for analytic control on the physics. Thus it will be applicable when the strain effects are small and the Kirigami structure doesn’t encounter any drastic modifications. However, as we will show these results are easily generalizable even in the presence of inhomogeneities and also open up new questions in non-homogeneous systems.
IV Multiscale methodology
We work in a ballistic regime where electronic transport can be estimated via the transmission properties of the network, which can in turn be estimated through a wavepacket evolution. However, as we will illustrate, it is useful both physically and numerically [46] to perform this analysis in three stages: (i) we first characterize the junction between adjoining blocks such that they are governed by a scattering matrix between the edge states. Any node in this network is composed of two incoming ($1,3$) and two outgoing channels ($2,4$) (see Fig. 2(a)). The scattering matrix $S$ is defined by two parameters $r$ and $t$ which are in general $\theta$ dependent
@@ IV Multiscale methodology
where $|r|^{2}+|t|^{2}=1$ [29]. An accurate estimation of $r,t$ at any node thus completely subsumes its local geometrical structure including any angular dependence, and in general $r,t$ can vary from node to node. (ii) After this first step, we show that the wave-packet dynamics can be modelled via effective non-Hermitian junction (ENHJ) where chiral edge states are replaced by one-dimensional non-Hermitian Hatano-Nelson (HN) models [47, 48] and the nodes are replaced by a four-site scattering junction as shown in Fig. 2(b). In a one-dimensional Hatano-Nelson (HN) model, fermions are allowed to hop in just one direction which can in turn model the edge states of a quantum Hall state. This uses the mapping of boundary theories of $d$ dimensional topological phases to $d-1$ dimensional non-Hermitian systems [49, 50, 51, 52] since this prevents any backscattering characteristic of a chiral system. Also, any non-homogeneity in the network can be effectively considered within the ENHJ formalism. (iii) In the final step of coarse-graining, one replaces each edge (side) with a single site, forming the median lattice of the network (see Fig. 2(c)). The non-Hermitian hopping amplitudes on this median lattice effectively model the time-evolution operator akin to the physics of Chalker-Coddington (CC) networks [27]. This analysis provides an analytic control to understand the phase diagram (see Fig. 1(c-d)) in a translationally invariant network. Moreover, this provides a controlled starting point for introducing inhomogeneities in the network. This multiscale approach also leads to numerically inexpensive simulations that can be easily generalized. The total size of simulations progressively changes from $2L^{2}L^{2}_{B}\rightarrow 4LL^{2}_{B}\rightarrow 4L^{2}_{B}$ under each step of coarse graining[46]. In the next section, we discuss the results of these methods as applied to our Kirigami network.
V Results
The discussions and results of the three methods as introduced in the earlier section are now discussed in detail. We first characterize the properties of nodes using wave packet dynamics.
V.1 Wave-packet dynamics